OpenStudy (anonymous):

A person can swim 0.65 m/s in still water. She heads directly south across a river 130 m wide and lands at a point 88 m [W] downstream. a)determine the velocity of the water relative to the ground. b)the swimmer's velocity relative to Earth. c) the direction she should swim to land at a point directly south of the starting point.

OpenStudy (anonymous):

time to cross = 130m/0.65 = 200s in 200s he moves 88m west, so speed = 88/200 = 0.44m/s west, relative to the ground Earth's speed? rotation speed = 1670 kilometers/hour, so 27.8km/s or 2780m/s, I suppose you would minus 0.65 from this? completely negligible. part c) using diagram, solve it for theta